OpenStudy (anonymous):

NEED HEELLPPPPPP What value of m makes the equation true? [–6 + (–13)] + 8 = –6 + (m + 8)

OpenStudy (anonymous):

@paki @sourwing Do you guys know this?

OpenStudy (mathmath333):

just solve the equation or observe the use of associative property here use

OpenStudy (triciaal):

get rid of parenthesis by distributing compare both sides of the equation and cancel same term -6-13 + 8 = -6 + m + 8 m = -13 yes use of associative property for addition the order of associating does not matter -6 -13 then + 8 = -19 + 8 = -11 -6 + m + 8 = m + 2 =-11 then m = -11 -2 =-13